Find the value of each of the following quantities:
C(8,4)=

C(12,7)=

C(12,8)=

C(11,1)=

C(10,9)=

C(8,6)=
1
Expert's answer
2020-11-02T20:42:31-0500

Formula:

"nC_k=\frac{n!}{k!(n-k)!}"


Solution:


C(8,4):

n=8;r=4

"8C_4=\frac{8!}{4!(8-4)!}" ="8C_4=\frac{8!}{4!(4)!}" =70




C(12,7): Here n=12, r= 7

"12C_7=\frac{12!}{7!(12-7)!}=792"



C(12, 8) n=12, r= 8

"= \frac{12!}{( 8! (12 - 8)! )}" =495



C(11,1) n=11, r= 1

"= \frac{11!}{( 1! (11 - 1)! )}" = 11



C(10,9) : n=10, r= 9

"= \frac{10!}{( 9! (10 - 9)! )}" = 10



C(8,6): n=8, r= 6

"= \frac{8!}{( 6! (8 - 6)! )}" = 28